# Topological consequences of null-geodesic refocusing and applications to $Z^{x}$ manifolds

### v2

@@ 4 Observer-Refocusing Spacetimes
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$ and let $\gamma$ be a timelike curve in $X$. Then the set $R_{p}^{\gamma}\coloneqq\{q\in\mathrm{im}(\gamma)\mid\exists v\in T_{p}X\text{ lightlike and future pointing and }t>0\text{ with }\alpha_{v}(t)=q\}$ is discrete.
@@
We can assume that $(X,g)$ is null-geodesically complete, else consider some conformally equivalent spacetime that is. Let $L$ be an embedded lightsphere at $p$. Assume that there is a sequence $(q_{n})_{n\in\mathbb{N}}$ with $q_{n}\in R_{p}^{\gamma}$ converging to some $q\in X$ with $q_{n}\neq q$ for all $n\in\mathbb{N}$. For each $n\in\mathbb{N}$, let $(v_{n},t_{n})\in L\times\mathbb{R}$ be such that $\alpha_{v_{n}}(t_{n})=q_{n}$. By Lemma 4.36, we know that $(t_{n})_{n\in\mathbb{N}}$ is bounded. Without loss of generality, assume that $(v_{n},t_{n})_{n\in\mathbb{N}}$ converges to some $(v,t)\in L\times\mathbb{R}$, else use compactness of $L$ and boundedness of $(t_{n})_{n\in\mathbb{N}}$ to go to a convergent subsequence.
Let $\beta:[0,1]\to L\times\mathbb{R}$ be a $\mathcal{C}^{1}$ curve with $\beta(0)=(v,t)$ and $\beta^{\prime}(0)\neq 0$ and such that for every $\varepsilon>0$, there are infinitely many $(v_{n},t_{n})\in\beta([0,1])$. For $s\in[0,1]$, let us write $\beta(s)=(v_{s},t_{s})$, abusing the notation slightly by giving new meaning to the terms to $v_{0},v_{1}$ and $t_{0},t_{1}$.
Consider the curve $\alpha(s)$ defined through $\alpha(s)\coloneqq\exp(t_{s}v_{s})=\alpha_{v_{s}}(t_{s})$. $\alpha$ is clearly $\mathcal{C}^{1}$. Further, for every $\varepsilon>0$, there are infinitely many intersections of $\alpha|_{[0,\varepsilon]}$ and $\gamma$. Let $t$ be such that $\alpha(0)=q=\gamma(t)$. Since both $\alpha$ and $\gamma$ are $\mathcal{C}^{1}$, the fact that $\alpha|_{[0,\varepsilon]}$ intersects $\gamma$ an infinite number of times for every $\varepsilon>0$ implies that $\alpha^{\prime}(0)$ and $\gamma^{\prime}(t)$ are linearly dependent. Since by construction $\alpha^{\prime}(0)\neq 0$, there is some $\lambda\neq 0$ with $\alpha^{\prime}(0)=\lambda\cdot\gamma^{\prime}(t)$. This is a contradiction to the fact that $\alpha^{\prime}(0)$ is not timelike due to Lemma 4.35. ∎
4.38 Corollary.
Let $(M,h)$ be a Riemannian manifold, $x,y\in M$. Then $R_{x}^{y}\coloneqq\{t>0\mid\text{ there exists a unit-speed geodesic }\gamma\text{ with }\gamma(0)=x\text{ and }\gamma(t)=y\}$ is a discrete set.
@@
Consider the associated spacetime $(X,g)$ of $(M,h)$ and set $p=(x,0)$ and $\gamma(t)=(y,t)$. Then $R_{x}^{y}=\gamma^{-1}(R_{p}^{\gamma})$ is discrete since $R_{p}^{\gamma}$ is by Proposition 4.37. ∎
4.39 Proposition.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$, let $\{q_{i}\}_{i\in I}\subset X$ be some countable set of points in $X$ and let $T>0$. If for each future-oriented lightlike geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in q_{i}$, then the Cauchy surface of $(X,g)$ is compact and has finite fundamental group.
Proof.
We can assume that $(X,g)$ is null-geodesically complete, else consider a conformally equivalent spacetime that is. For each $i\in I$, consider $K_{i}\coloneqq\{v\in L\mid\exists t\in(0,T]\mid\alpha_{v}(t)=q_{i}\}$. Let $v\in\mathrm{int}_{L}(K_{i})$. Then there is a (non-constant) null-geodesic variation through $v$, with each null-geodesic going through $p$ and $q_{i}$. This means that $q_{i}$ is conjugate to $p$ along $\alpha_{v}|_{[0,T]}$. Hence $\alpha_{v}(t)\in I^{+}(p)$ for all $t>T$. Let $A=L\setminus\bigcup_{i\in I}\partial_{L}(K_{i})$, then $A$ is dense by the Baire category theorem. Further, if $v\in A$, then there is some $i\in I$ with $v\in K_{i}\setminus\partial_{L}(K_{i})=\mathrm{int}_{L}(K_{i})$, and hence $\alpha_{v}(t)\in I^{+}(p)$ for all $t\geq T$. Since $A$ is dense, $X$ has compact Cauchy surface.
To show that the Cauchy surface of $(X,g)$ has finite fundamental group, we proceed as follows. Let $\rho:(\tilde{X},\tilde{g})\to(X,g)$ be the semi-Riemannian universal cover of $X$. Let $\tilde{p}$ be any point in $\rho^{-1}(p)$ and let $\tilde{L}$ be the lightsphere at $p$ given by $\tilde{L}=(d\rho)^{-1}(L)$. Note that the set $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Let $\tilde{v}\in\tilde{L}$ and $v=d\rho(\tilde{v})\in L$. Then there is some $t\in(0,T]$ and $i\in I$ with $\alpha_{v}(t)=q_{i}$. Lifting $\alpha_{v}$ to $\tilde{X}$, we get that $\alpha_{\tilde{v}}(t)\in\rho^{-1}(q_{i})\in\rho^{-1}(\{q_{i}\}_{i\in I}\})$. Note that $\rho^{-1}(\{q_{i}\}_{i\in I}\})$ is countable. Through application of the first part of this theorem, we can conclude that $\tilde{X}$ has compact Cauchy surface as well. Now pick a Cauchy surface $S\subset X$ and let $\tilde{S}=\rho^{-1}(S)$. Then $\tilde{S}$ is a Cauchy surface of $\tilde{X}$ (see the proof of [15, Theorem 11.5]) and hence compact. Further, $\tilde{S}$ is simply connected (as $\tilde{X}$ is simply connected and diffeomorphic to $\tilde{S}\times\mathbb{R}$), and $\rho:\tilde{S}\to S$ is a covering map. Pick any $x\in\tilde{S}$, then
#$\#\pi_{1}(S,x)=\#\rho^{-1}(x),$
which is finite since $\tilde{S}$ is compact. ∎
4.40 Theorem.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$, let $\{\gamma_{i}\}_{i\in I}$ be some countable family of timelike curves and let $T>0$. If for each future-oriented lightlike geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in\mathrm{im}(\gamma_{i})$, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
Proof.
For each $i\in I$ the set $R_{p}^{\gamma_{i}}$ is discrete, hence countable. The claim follows immediately from applying the Proposition 4.39 to the countable set $\bigcup_{i\in I}R_{p}^{\gamma_{i}}$. ∎
4.41 Proposition.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ such that the domain of $\gamma$ is compact, then $(X,g)$ has compact Cauchy surface and finite fundamental group.
Proof.
Using similar arguments as in Lemma 4.36 compactness of $\mathrm{im}(\gamma)$ ensures that there is some lightsphere $L$ at $p$ and some $T>0$ such that for all $v\in L$ there is some $t\in(0,T]$ with $\gamma(t)\in\mathrm{im}(\gamma)$. Now we can use Theorem 4.40
4.42 Corollary.
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ such that all unit-speed geodesics starting from $x$ reach $y$ in uniformly bounded time. Then $M$ is compact with finite fundamental group.
Proof.
Let $T>0$ be the uniform bound on the time it takes unit-speed geodesics starting at $x$ to reach $y$. Consider the associated spacetime $(X,g)$ to $(M,h)$. Defining $p=(x,0)$ and $\gamma:[0,T]\to X$, $\gamma(t)=(y,t)$, the result then follows from Proposition 4.41. ∎
4.43 Corollary.
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension $2$ or $3$, let $x\in M$ and assume that all unit-speed geodesic starting at $x$ reach $y$ in uniformly bounded time. Then there exists a Riemannian metric $\tilde{h}$ on $M$ such that $(M,\tilde{h})$ is a $Y^{x}_{l}$ manifold.
Proof.
In this case $M$ is compact and has finite fundamental group. If $\dim(M)=3$, then by the Thurston Elliptization conjecture [30] proved by Perelman [26], [27], [28], $M$ is a quotient of the sphere $S^{3}$ (with standard metric) by a finite group of isometries. The induced metric on $M$ will make it into a $Y^{x}_{2\pi}$ manifold. Similar arguments work for $\dim(M)=2$. ∎
4.44 Remark.
If we drop the requirement of a uniform bound from Theorem 4.40 we can still show that there is an open, dense subset $A\subset L$ such that all null-geodesics $\alpha$ starting at $p$ with initial velocity in $A$ eventually enter $I^{+}(p)$, but this is not enough to conclude that we have a compact Cauchy surface.
4.45 Remark.
It is unknown whether there are $Z^{x}$ manifolds where the time-of-return of unit-speed geodesics is not uniformly bounded [9, Question 7.71]. Similarly it is unknown whether there are any globally hyperbolic observer-refocusing spacetimes where the timelike curve cannot be taken to be compact.
5 The Analytic Case
5.46 Lemma.
Let $(X,g)$ be a globally hyperbolic spacetime with $g$ being an analytic metric. Then there exists an analytic, surjective temporal function $\mathcal{T}:X\to\mathbb{R}$ such that every level set of $\mathcal{T}$ is a spacelike Cauchy hypersurface.
Proof.
Since $X$ is globally hyperbolic, Bernal and Sanchez proved in [4], [5] that there exists an isometry between $(X,g)$ and $(S\times\mathbb{R},\hat{g})$, where $S$ is a Cauchy surface of $X$ and $\hat{g}$ is a metric of the form $\hat{g}=h-Adt^{2}$ for some smooth function $A:S\times\mathbb{R}\to\mathbb{R}^{+}$ and so that $h$ is a symmetric tensor which restricts to a Riemannian metric on every hypersurface $S\times\{t\}$ and with $h(\partial_{t},\cdot)=0$. Let $\hat{\mathcal{T}}:S\times\mathbb{R}\to\mathbb{R}$ be the projection onto the time-coordinate, then $\hat{\mathcal{T}}$ is an analytic, surjective temporal function with all level sets spacelike Cauchy surfaces. By a theorem of Myers and Steenrod [23], isometries are always at least as regular as the metric. Therefore the isometry $\varphi:(X,g)\to(S\times\mathbb{R},\hat{g})$ is analytic. The temporal function $\mathcal{T}\coloneqq\hat{\mathcal{T}}\circ\phi:X\to\mathbb{R}$ is then analytic as the composition of two analytic functions. ∎
5.47 Theorem.
Let $(X,g)$ be a globally hyperbolic, observer-refocusing spacetime of dimension at least $3$, with $g$ an analytic metric. Then $(X,g)$ is strongly refocusing.
Proof.
@@ 5 The Analytic Case
Let $p\in X$ and let $\gamma$ be a timelike curve in $X$ so that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. Let $L\subset T_{p}X$ be some lightsphere at $p$ which is also an analytic submanifold of $T_{p}X$, i.e. a lightsphere such that the inclusion $\iota:L\to T_{p}X$ is analytic. Let $D\subset L\times\mathbb{R}$ be the maximal domain of the function $(t,v)\mapsto\exp_{p}(tv)$. Note that the map $H:D\to T_{p}X$ defined through $H(v,t)=t\cdot\iota(v)$ is analytic. Let $\mathcal{T}$ be a surjective, analytic temporal function such that each level set of $\mathcal{T}$ is a Cauchy surface and consider the map
@@
For each $q\in R_{p}^{\gamma}$ consider the set $K_{q}\coloneqq\{v\in L\mid\exists t>0\text{ with }\alpha_{v}(t)=q\}$. Each $K_{q}$ is a closed set and $L=\bigcup_{q\in R_{p}^{\gamma}}K_{q}$ is a countable union of these sets. By the Baire category theorem, there has to be some $q\in R_{p}^{\gamma}$ such that $U\coloneqq\mathrm{int}_{L}(K_{q})$ is non-empty. Let $T=\mathcal{T}(q)$. Then, for every $v\in L$, there is a unique $t\in\mathbb{R}$ with $(v,t)\in D$ and $\phi(v,t)=T.$ By the analytic implicit function theorem, there is an analytic map $f:L\to\mathbb{R}$ with $(v,f(v))\in D$ and $\phi(v,f(v))=T$ for all $v\in L$. Now consider the map
@@
which is analytic since the maps $\exp_{p}$, $H$, and $f$ are analytic. For $v\in U$ we must have $\phi(v)=q$, so $\psi$ is constant on $U$. Since $\psi$ is analytic on all of $L$, $\psi$ must be constant on $L$. ∎
@@
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with $h$ an analytic metric, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
@@ 6 Analogous conjecture in the Contact case
The results of Chernov and Nemirovski [13], the results of Frauenfelder, Labrousse and Schlenk [16, Theorem 1.11 and 1.13], and our results about Riemannian $Z^{(x,y)}$ manifolds and observer-refocusing spacetimes in this paper, suggest the following conjectures.
@@
Let $M$ be manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\alpha$ be any contact form on $(S^{\ast}M,\xi)$, i.e. $\xi=\ker(\alpha)$, let $\varphi_{\alpha}^{t}$ be the Reeb flow of $\alpha$ and let $x,y\in M$. Assume that for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
This contact-geometric version generalizes the Riemannian $Z^{(x,y)}$ version as follows: Let $(M,g)$ be a Riemannian manifold, let $\psi:TM\to T^{\ast}M$ be the isomorphism defined through $v\mapsto g(v,\cdot)$, then $\psi$ induces a bundle-isomorphism (which we will also denote by $\psi$) between the unit-tangent bundle $U_{g}M$ and the spherical cotangent bundle $S^{\ast}M$. We can define a $1$-form $\alpha$ on $S^{\ast}M$ via $\alpha_{v}(w)=g(\psi^{-1}(v),d\pi(w))$ where $\pi:S^{\ast}M\to M$ is the natural projection. The Reeb flow of $\alpha$ on $S^{\ast}M$ corresponds to the co-geodesic flow of $g$ on $S^{\ast}M$. Conjecture 6.49 generalizes the conjecture that $Z^{x}$ manifolds of dimension at least $2$ have finite fundamental group. We have proven the compactness and the fundamental group aspect of Conjecture 6.50 whenever the contact form $\alpha$ comes from a Riemannian metric. The conjecture states that this is really a fact about Reeb flows on $(S^{\ast}M,\xi)$ and not just about geodesics on $M$. One may also ask other questions, for example if the set $R_{x}^{\alpha}\coloneqq\{t>0\mid\exists v\in S^{\ast}_{x}M\text{ such that }\varphi_{\alpha}^{t}(x)\in S^{\ast}_{x}M\}$ is discrete for all contact forms $\alpha$ on $(S^{\ast}M,\xi)$. If it is always discrete, then Conjecture 6.49 holds for analytic $1$-forms $\alpha$ by the analyticity of the Reeb flow and by the analyticity of $\pi:S^{\ast}M\to M$ because arguments as in Theorem 5.47 can reduce the problem to the case already solved by Frauenfelder, Labrousse and Schlenk.
@@
Let $M$ be manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\phi_{t}$ be any positive Legendrian isotopy of $(S^{\ast}M,\xi)$ with $\phi_{0}=\mathrm{id}$ and let $x,y\in M$. If for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\phi_{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
Let $M$, $\xi$, $\alpha$ and $\phi_{t}$ be as in Conjecture 6.52. Let $x,y\in M$. Assume there is some $T>0$ such that for all $v\in S^{\ast}_{x}M$ there is some $t\in(0,T]$ with $\varphi_{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.

### v4

@@ 5 The Analytic Case
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ in $X$. Let $L$ be a lightsphere at $p$. Then there exists a non-empty open subset $V\subseteq L$ and a point $q\in\mathrm{im}(\gamma)$ such that for all $v\in V$ there is some $t>0$ with $\alpha_{v}(t)=q$
@@
As shown in the proof of Theorem 4.40, there is an open, dense subset $L^{\prime}\subseteq L$ such that for all $v\in L^{\prime}$, $p$ has some conjugate point $q$ in the intersection of $\alpha_{v}$ and $\gamma$. We define $f:L^{\prime}\to\mathbb{R}^{+}$ through $\exp(f(v)v)=\text{first conjugate point }q\in\mathrm{im}(\gamma)\text{ along }\alpha_{v}.$ It is easy to see $f$ is lower semi-continuous. By Lemma 5.49, $v\mapsto\exp(f(v)v)$ is constant on some non-empty open set $V\subseteq L^{\prime}$. In particular, there is some $q\in\mathrm{im}(\gamma)$ such that for all $v\in V$ there is some $t=f(v)>0$ such that $\alpha_{v}(t)=\exp(tv)=q$. ∎
5.51 Theorem.
Let $(X,g)$ be a globally hyperbolic, observer-refocusing spacetime of dimension at least $3$ with an analytic metric $g$. Then $(X,g)$ is strongly refocusing.
@@
By [33, Theorem 2.10], there exists an analytic temporal function $\mathcal{T}:X\to\mathbb{R}$ whose level sets are Cauchy surfaces. Let $p\in X$ and let $\gamma$ be a timelike curve in $X$ so that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. Let $L\subset T_{p}X$ be some lightsphere at $p$ which is also an analytic submanifold of $T_{p}X$, i.e. a lightsphere such that the inclusion $\iota:L\to T_{p}X$ is analytic. Let $D\subset L\times\mathbb{R}$ be the maximal domain of the function $(v,t)\mapsto\exp_{p}(tv)$. Note that the map $H:D\to T_{p}X$ defined through $H(v,t)=t\cdot\iota(v)$ is analytic. Consider the map
@@
By Proposition 5.50 there is some non-empty open set $V\subseteq L$ and a point $q\in\mathrm{im}(\gamma)$ such that for all $v\in V$ there is some $t>0$ with $\alpha_{v}(t)=q$. Let $T=\mathcal{T}(q)$. Then, for every $v\in L$, there is a unique $t\in\mathbb{R}$ with $(v,t)\in D$ and $\phi(v,t)=T.$ By the analytic implicit function theorem, there is an analytic map $f:L\to\mathbb{R}$ with $(v,f(v))\in D$ and $\phi(v,f(v))=T$ for all $v\in L$. Now consider the map
@@
which is analytic since the maps $\exp_{p}$, $H$, and $f$ are analytic. Note that for $v\in V$ we must have $\psi(v)=q$, so $\psi$ is constant on $V$. Since $L$ is topologically a sphere of dimension $n-2\geq 1$, $L$ is connected. Since $\psi$ is analytic on all of $L$ and constant on the non-empty open set $V$, $\psi$ is constant on all of $L$. ∎
5.52 Remark.
Note that we did not need to assume that $(X,g)$ is observer-refocusing with respect to an analytic timelike curve $\gamma$. We only need analyticity of the metric $g$.
@@
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ with an analytic metric $h$, then $(M,h)$ is a $Y^{(x,y)}_{l}$ manifold for some $l>0$.
Proof.
This result follows from the proof of Theorem 5.51 by considering the associated spacetime $(X,g)$ to the Riemannian manifold $(M,h)$, which will be analytic since $(M,h)$ was analytic, defining $p=(x,0)$, and the standard observer $\gamma(t)=(y,t)$. ∎
@@ 6 Analogous conjecture in the Contact case
The results of Chernov and Nemirovski [14], the results of Frauenfelder, Labrousse and Schlenk [18, Theorem 1.11 and 1.13], and our results about Riemannian $Z^{(x,y)}$ manifolds and observer-refocusing spacetimes in this paper, suggest the following conjectures.
@@
Let $M$ be a manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\alpha$ be any contact form on $(S^{\ast}M,\xi)$, i.e. $\xi=\ker(\alpha)$, let $\varphi_{\alpha}^{t}$ be the Reeb flow of $\alpha$ and let $x,y\in M$. Assume that for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\varphi_{\alpha}^{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
This contact-geometric version generalizes the Riemannian $Z^{(x,y)}$ version as follows: Let $(M,g)$ be a Riemannian manifold, let $\psi:TM\to T^{\ast}M$ be the isomorphism defined through $v\mapsto g(v,\cdot)$. Then $\psi$ induces a bundle-isomorphism (which we will also denote by $\psi$) between the unit-tangent bundle $U_{g}M$ and the spherical cotangent bundle $S^{\ast}M$. We can define a $1$-form $\alpha$ on $S^{\ast}M$ via $\alpha_{v}(w)=g(\psi^{-1}(v),d\pi(w))$ where $\pi:S^{\ast}M\to M$ is the natural projection. The Reeb flow of $\alpha$ on $S^{\ast}M$ corresponds to the co-geodesic flow of $g$ on $S^{\ast}M$. Conjecture 6.54 generalizes the conjecture that $Z^{x}$ manifolds of dimension at least $2$ have finite fundamental group. We have proven the compactness and the fundamental group aspect of Conjecture 6.55 whenever the contact form $\alpha$ comes from a Riemannian metric. The conjecture states that this is really a fact about Reeb flows on $(S^{\ast}M,\xi)$ and not just about geodesics on $M$.
@@
Let $M$ be a manifold of dimension at least $2$, let $(S^{\ast}M,\xi)$ be the spherical cotangent bundle on $M$ equipped with the standard contact structure. Let $\phi_{t}$ be any positive Legendrian isotopy of $(S^{\ast}M,\xi)$ with $\phi_{0}=\mathrm{id}$ and let $x,y\in M$. Assume that for all $v\in S^{\ast}_{x}M$ there is some $t>0$ with $\phi_{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite, and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.
@@
Let $M$, $\xi$, and $\phi_{t}$ be as in Conjecture 6.57. Let $x,y\in M$. Assume there is some $T>0$ such that for all $v\in S^{\ast}_{x}M$ there is some $t\in(0,T]$ with $\phi_{t}(v)\in S^{\ast}_{y}M$. Then $M$ is compact, the fundamental group of $M$ is finite and the integral cohomology ring of the universal cover of $M$ is the one of a CROSS.